Circular No. 7403

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/2000 ET_{90} (KOWAL-MRKOS) MPS 11479 contains observations on Mar. 9 and 13 by LINEAR of

an apparently asteroidal object presumed to have a moderately eccentric orbit in the inner part of the main belt. Linkage by G. V. Williams to LINEAR observations on Apr. 4 and 8 demonstrated the cometary nature of the orbit, and the object was placed in The NEO Confirmation Page. Isolated observations from LINEAR on Feb. 7, from the Catalina Sky Survey on Mar. 1 (when observer T. B. Spahr had in fact drawn attention to the object’s “slowish” motion) and from LONEOS on Apr. 2 were then also linked. Neither these observers nor those responding to the Confirmation Page made a definite remark about the object’s cometary appearance, even in response to specific enquiries from the Central Bureau (although strong moonlight has recently been a factor). Following a suspicion by the undersigned and an independent suggestion by C. W. Hergenrother, 2000 ET_{90} has been definitively identified with comet D/1984 H1 = 1984 JD (Kowal-Mrkos) = 1984n (IAUC 3988, 4001) = 1984 X, for which current predictions (ICQ Comet Handbook for 2000, p. H87; OAA Comet Handbook for 2000, p. 37) require correction by ΔT ∼ −125 days. The comet passed only 0.16 AU from Jupiter in Mar. 1989. The rediscovery position and orbital elements (from 47 observations 1984–2000, mean residual 0′′.7) follow. There was an unobserved return with T = 1991 Aug. 2. The hitherto unpublished astrometry, residuals and an ephemeris are given on MPEC 2000-G52.

2000 UT  α_{2000}  δ_{2000}  m_{2}
Mar. 9.30345  11^{h}38^{m}37.^{s}03  −5^{o}21′29′′.0  18.9

Epoch = 1984 May 20.0 TT

T = 1984 June 8.3272 TT  ω = 338.^{o}3150
e = 0.484237  Ω = 249.3254 }2000.0
q = 1.949042 AU  i = 2.9538

a = 3.778947 AU n^{o} = 0.1341677 P = 7.346 years

Epoch = 2000 June 25.0 TT

T = 2000 July 1.8035 TT  ω = 320.^{o}5886
e = 0.409022  Ω = 245.5014 }2000.0
q = 2.546791 AU  i = 4.6837

a = 4.309452 AU n^{o} = 0.1101720 P = 8.946 years

2000 April 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Brian G. Marsden